An isolated sphere, A has a radius 18cm and carries a charge of 4.0x10^-7C

(a) calculate the potential of the sphere

(b) another isolated sphere, B has a radius 9cm and is ...

(c) if sphere A and B are now connected with a wire, what is the relationship between the potential of A and B

(d) will charges flow from A and B or from B to A when the spheres are connected with wire?

(e) calculate the charge transferred from one sphere to the other in number d

(f) calculate also the final charge in each sphere, and

(g) calculate the final potential of each sphere


1
Expert's answer
2021-03-23T11:12:34-0400

a)


"V=9\cdot10^9\frac{4\cdot10^{-7}}{0.18}=20000\ V"

b)


"V=9\cdot10^9\frac{4\cdot10^{-7}}{0.09}=40000\ V"

c)


"V_A=V_B"

d) From B to A.

e)


"\frac{4\cdot10^{-7}+q}{0.18}=\frac{4\cdot10^{-7}-q}{0.09}\\q=1.33\cdot10^{-7}\ C"

f)


"q_B=5.33\cdot10^{-7}\ C\\q_A=2.67\cdot10^{-7}\ C"

g)


"V'=9\cdot10^9\frac{5.33\cdot10^{-7}}{0.18}=26650\ V"